OpenStudy (anonymous):

Sean is arranging 4 decorations on fireplace mantel. How many arrangements are possible?

OpenStudy (anonymous):

\(4!\)

OpenStudy (anonymous):

in other words \(4\times 3\times 2\)

OpenStudy (anonymous):

how did you get 3x2

OpenStudy (anonymous):

4! = 4x3x2x1

OpenStudy (anonymous):

ikr^ idont get it lol

OpenStudy (anonymous):

you have 4 decorations, lets call them A, B, C and D and you are going to arrange them in some order on your mantel. lets make the slots ____ ____ ____ _____ so one possibility could be A C B D ____ ____ ____ ____

OpenStudy (anonymous):

question is how many ways are there to do this? first pick one to go in the first slot you have 4 choices now pick one to go in the second slot you have 3 choices (since one has been chosen) then pick one to go in the third slot you have 2 choices (since two have been chosen) now you have no choices for the last slot since you have placed the other 3 already by the counting principle, the number of ways to do all these together is the product of the number of ways to do them individually, so it is \[4\times 3\times 2=24\]

OpenStudy (anonymous):

so the answer is 24? not 4

OpenStudy (anonymous):

the answer is 24 and \(4!\) does not mean really really it is 4, \(4!\) is short notation for \[4\times 3\times 2\] and is called 4 "factorial"

OpenStudy (anonymous):

for example \[6!=6\times 5\times 4\times 3\times 2\]

OpenStudy (anonymous):

oh my gosh . I'm only in 7th grade and I don't understand this lol

OpenStudy (anonymous):

so what should I put on the answer slot lol

OpenStudy (anonymous):

yes of course if you have not seen it before you do not understand it, but i hope it is clearer now counting principle says if there are 5 ways to do one thing and 3 ways to do another, then there are 15 ways to do them together more generally if there are n ways to do one thing and m ways to do another then there are \(m\times n \) ways to do the two tasks together

OpenStudy (anonymous):

the answer is 24 the equation is 4! which is = to 4x3x2x1 and that multiplies out to 24